OpenStudy (anonymous):

5. Is (15, 5) a solution to the equation 2x = 5y + 5?

OpenStudy (anonymous):

Well, 15=x and 5=y Sub them in and check!

OpenStudy (anonymous):

Can you please show your work so i know how you did it please

OpenStudy (anonymous):

Um... there isn't a lot of work. 2(15)=5(5)+5 30=25+5 30=30

OpenStudy (anonymous):

oh really awe thank u